Assume that you have been hired as a consultant by CGT, a major producer of chemicals and plastics, the company has long term debt/equity ratio of 0.5. The stock is currently selling for $15.25 per share, and its $1,000 par value, 20-year, 7.25% bonds with semiannual payments are selling for $875.00. The beta is 1.25, the yield on a 6-month T-bill is 3.50%, and the yield on a 20-year Treasury bond is 5.50%. The required return on the stock market is 11.50%, but the market has had an average annual return of 14.50%. The firm's tax rate is 40%.

What is the best estimate of the after-tax cost of debt?

Based on the CAPM, what is the firm's cost of common stock?

Which of the following is the best estimate for the weight of debt for use in calculating the firm’s WACC?

What is the best estimate of the firm's WACC?

PLEASE NO EXCEL
Answer #1

Cost of debt can be calculated using I/Y function on a calculator

N = 20 x 2 = 40, PMT = 7.25% x1000 / 2 = 36.25, PV = -875, FV = 1000 => Compute I/Y = 4.28% (semi-annual)

After-tax cost of debt, rd = 4.28% x 2 x (1 - 40%) = 5.14%

Using CAPM, Cost of equity, re = Rf + beta x (Rm - Rf) = 3.5% + 1.25 x (11.50% - 3.5%) = 13.50%

Weight of debt, wd = D/E / (1 + D/E) = 0.5 / 1.5 = 33.33%, Weight of equity, we = 1 - 33.33% = 66.67%

WACC = wd x rd + we x re = 33.33% x 5.14% + 66.67% x 13.50% = 10.71%